In a hot dog eating contest, Jonah eats 12 hot dogs in 2 minutes. What is Jonah's rate of hot dog consumption? Express your answer in lowest terms.
Answer by darknez25(6)

No. of hot dogs eaten in 2 mins = 12
No. of Hot dogs eaten in 1 mins = 12/2
No. of hot dogs eaten in 60 secs = 12/2
No. of hot dogs eaten in 10 secs = 12/(2 X 6)
No. of hot dogs eaten in 10 secs = 12/12
No. of hot dogs eaten in 10 secs = 1
Therefore the rate of consumption is 1 Hot dog per 10 seconds.
ie 1 hotdog / 10 secs